TERRI'S WEEKLY INCOME IS A $100 DOLLARS LESS THAN DOULBE GARY'S WEEKLY IONCOME. IF TERRI MAKE'S $448 A WEEK, WHAT DOES GARY MAKE?

AND I HAVE A NOTHER QUESTION
PAUL MAKES $25 A WEEK LEES THAN THE SUM OF WHAT AVERY AND CASEY MAKES TOGETHER MAKE. CASEY'S WEEKLY INCOME WOULD BE TRIPLE STEVE'S IF CASEY MAKES $50 MORE A WEEK. PAUL MAKES $342 A WEEK. HOW MUCH D OES AVERY MAKE?

Answer by sdmmadam@yahoo.com(530)

TERRI'S WEEKLY INCOME IS A $100 DOLLARS LESS THAN DOULBE GARY'S WEEKLY IONCOME. IF TERRI MAKE'S $448 A WEEK, WHAT DOES GARY MAKE?
Let Gary make G dollars a week.
Given that Terry's weekly income is $448
Given that TERRI'S WEEKLY INCOME IS A $100 DOLLARS
LESS THAN DOULBE GARY'S WEEKLY IONCOME
That is 448 is $100 < 2G
That is 448 = 2G - 100
That is (448+100) = 2G
548 = 2G
Therefore G = 548/2 = 274
Therefore Gary's weekly income is 274 dollars
Verification: Terry' should be 100 less than twice Gary's
And 448 is of course 100 less than (2X274)= 548

PAUL MAKES $25 A WEEK LEES THAN THE SUM OF WHAT AVERY AND CASEY MAKES TOGETHER MAKE. CASEY'S WEEKLY INCOME WOULD BE TRIPLE STEVE'S IF CASEY MAKES $50 MORE A WEEK. PAUL MAKES $342 A WEEK. HOW MUCH D OES AVERY MAKE?
Here in this problem information about Steve's weekly income is left out
Let Avery make A dollars.
Let Casey's weekly income be C $ and
Let Steve's weekly income be S $
We shall assume that S is known

Paul makes $342 a week
Paul makes $25 < (Avery+ Casey)
342 is $25 < (A+C)
That is 342 = (A+C)-25 ----(1)
342+25 = (A+C)
That is (A+C) = 367 ----(2)
casey's weekly income
= Three times Steve's weekly income if Casey makes $50 more
C +50 = 3S
That is C = 3S-50 ----(3)
Putting (3) in (2)
A+C = 367
A+(3S-50) =367
A = 367+50-3S
That is A = 417 - 3S
Answer: Avery makes (417-3S) dollars where S is Steve's weekly income.